OpenStudy (anonymous):

Which is the slope of the line with equation 4x 2y = 4? (Points : 3) Option A: -2 Option B: -1\2 Option C: 1\2 Option D: 2

OpenStudy (anonymous):

4x-2y=4?

OpenStudy (anonymous):

-4-2y=4, sorry

OpenStudy (anonymous):

either way.. solve this for y to get it in y=mx+b form so you can identify the slope. the slope will be the coefficient of x when you get it into that form.

OpenStudy (anonymous):

-4x-2y=4

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

how you gonna get that y by itself?

OpenStudy (anonymous):

I have no clue...

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

so we have to add it to both sides. right? since it is called an equation? it has to be equal? equal means equation.

OpenStudy (anonymous):

so lets add 4x to both sides.

OpenStudy (anonymous):

what are u left with now.

OpenStudy (anonymous):

+4x-4x-2y=4+4x right?

OpenStudy (anonymous):

+4x-4x=?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so you are left with -2y=4+4x.. now, since -2y represents -2 times y... how do you undo multiplication?

OpenStudy (anonymous):

you gone?